# Rotary Masked Autoencoders are Versatile Learners

### v1

@@ 5.4 Interpolation
We evaluate RoMAE on three interpolation tasks with increasing dimensionality and sampling irregularity. (i) Spiral: A 2D synthetic benchmark of 300 noisy Archimedean spirals as in Ref. [12]; (ii) Synthetic: The 50-step univariate task from Ref. [46] and (iii) PhysioNet: 48-hour ICU records containing 41 clinical variables [49]. For (i), each spiral is discretized into 75 evenly-spaced time steps. To create irregular time-series data, time points are randomly selected from the first half of each spiral, which are used for interpolation. For (ii), the interpolation task is between a random subsample including between 3 and 10 points per trajectory. For (iii), we follow the 50% masking protocol of [35]: half of the time, rows with at least one observation are hidden and must be reconstructed from context. We provide additional experimental details in the Appendix, respectively 16, C.7, and C.8. All experiments are conducted to ensure a direct comparison with the results of Transformer [59], LatentODE [9], ContiFormer [12] for (i) as reported by [12], and for HeTVAE for (ii) and (iii) as reported by [35]. Across scales, our RoMAE-tiny configuration consistently scores competitively with respect to benchmarked MSE/RMSE, as seen in Table. 6.
Table 6: Results for the interpolation experiments discussed in Section 5.4.
@@
HeTVAE [35] $\mathbf{0.562\pm 0.022}$
RoMAE $0.570\pm 0.014$
@@
RoMAE $\mathbf{0.0183\pm 0.007}$
@@ 6 Discussion
Table 6 shows that one tiny/small RoMAE model, pre-trained once with a generic masked-autoencoder objective and no task-specific architectural tuning, matches or surpasses specialised baselines across three increasingly difficult interpolation datasets. On the 2-D Spiral benchmark, we improve by an order of magnitude over the previous best result by ContiFormer, which we attribute to MAE tubelet-masking enforcing long-range reasoning. For the 50-step Synthetic task we obtain $0.233\pm 0.007$ MSE, comparable to HeTVAE’s $0.223\pm 0.070$ but with markedly tighter standard deviation (five seeds). On the irregular, 41-channel PhysioNet-2012 ICU data we achieve $0.570\pm 0.014$ MSE, within half a standard deviation of HeTVAE’s heteroscedastic decoder. We have verified by retraining HeTVAE on the the PhysioNet interpolation task that it indeed excels on densely sampled variables, such as heart-rate, whereas RoMAE maintains more balanced interpolation across the sparsely observed channels; consequently, while the aggregate MSE over all the features is similar, RoMAE delivers a simpler, single-stage model without task-specific adaptations.
@@
Limitations: RoPE in RoMAE has some additional computational overhead if the positions are different with each forward pass, e.g., with any continuous irregular time-series. If positions stay constant, however, as in images, the overhead becomes negligible. RoMAE is not well suited for very long sequences, as it uses standard Attention which has $O(n^{2})$ memory complexity with regards to sequence length. Lastly, RoMAE’s ability to perform on extrapolation tasks is limited, as discussed in Sec. A.5.
@@
7 Acknowledgments
RT acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. RT and MV are partially supported by the Fondazione ICSC, Spoke 3 “Astrophysics and Cosmos Observations”, Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19)” - Next Generation EU (NGEU). We also acknowledge the use of computational resources provided by the Italian AREA Science Park supercomputing platform ORFEO in Trieste. GC is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Postdoctoral Fellowship Programme, SMASH co-funded under the grant agreement No. 101081355. The operation (SMASH project) is co-funded by the Republic of Slovenia and the European Union from the European Regional Development Fund.
@@ 8 Conclusion
Appendix A Additional Experiments, Discussion, and Model Details
A.1 Architectural and Normalization Details
@@
$\text{RMS}(\mathbf{a})=\sqrt{\frac{1}{n}\sum_{i=1}^{n}a_{i}^{2}}\,,$ (3)
where $\mathbf{a}$ is a sequence of token embeddings. Notably, RMSNorm does not centre the input $\mathbf{a}$ like LayerNorm [2] does. Re-centring has been shown not to be necessary and removing it saves compute.
@@
A.2 Model Sizes
Table 7: All RoMAE model sizes.
Parameter tiny-shallow tiny small base
$d_{\text{model}}$ $180$ $180$ $432$ $720$
$N_{\text{head}}$ $3$ $3$ $6$ $12$
Depth $2$ $12$ $12$ $12$
dim feed-forward $720$ $720$ $1728$ $2880$
A.3 Additional Absolute Position Reconstruction Results
@@
Here we provide an additional experiment showing how RoMAE is able to to reconstruct absolute position across a wide range of values. To conduct the experiment, we pass only one token into RoMAE, giving it a random position drawn from a uniform distribution $\mathcal{U}_{[a,\ b]}$, training the model to predict this position. We also pass in the [CLS] token. The experiment is conducted over two domains; (0, 100)0100(0,\ 100)( 0 , 100 ) and (0, 1000)01000(0,\ 1000)( 0 , 1000 ). Hyperparameters and training details are discussed in Section C.5. After training, we evaluate how well the model performs across the range of values it was trained on. Results are plotted in Figure 2.
@@
Discussion: We find that the model is generally able to reconstruct all position values within the two domains tested, except when the position is close to the edges of the domain. This effect occurs already when the domain is relatively small and worsens as it becomes larger. We also find that the model performs better overall on the smaller domain. These results provide empirical support for Proposition 4.2, and show how the model is able to learn to reconstruct absolute position across a large domain. That the loss grows as the position nears the edges of the domain shows that the model does not find solutions that generalize to out-of-distribution positions.
A.4 Audio benchmark
We chose to test RoMAE’s ability to classify audio files, after a self-supervised pre-training on unlabeled audio datasets, inspired by the SSAST pretraining strategy [21]. SSAST is the first transformer-based model that introduces a self-supervised pretraining strategy, supporting arbitrary patch size, for the audio representation learning.
In order to apply our model to audio datasets, we first transform the audio waveforms to mel spectrograms – a spectrogram where the frequencies are converted to the mel scale. First, the input audio waveform of length $t$ seconds is converted into a sequence of 128-dimensional log-Mel filterbank [^4] Mel Filterbank transform computes weighted averages of bins to provide spectral power estimates on a logarithmic frequency scale, which is more affine to human hearing resolution. (fbank) features computed with a 25ms Hamming window every 10ms. This results in a 128 × 100 $t$ spectrogram. For the pretraining step, we followed the patchification strategy adopted in [21] and we split the spectrogram into a sequence of $N$ $16\times 16$ patches, with an overlap of 6 in both time and frequency dimension, where $N=12(100t-16)/10$ is the number of patches and the effective input sequence length for the model. In Table 8 we report the hyperparameters configuration used for the pretraining phase.
Table 8: Pretraining hyperparameters for audio benchmark.
Hyperparameter Value
Model size Small
Optimizer AdamW
AdamW betas $\beta_{1}=0.9,\ \beta_{2}=0.999$
Weight Decay 0.05
Base LR $8\times 10^{-4}$
Batch size 24
Epochs 50-100
Linear LR warmup steps 500
LR schedule Cosine
Dropout 0
Stochastic depth $\lambda$ 0
Precision FP32
For the benchmark we used the ESC-50 dataset [39], consisting of 2000 5-second environmental audio recordings classified into 50 classes. The current best results on ESC-50 are accuracies of 86.5 and 94.7 obtained with supervised training (on AudioSet-2M) respectively by SOTA-S and SOTA-P models. The SSAST model, which is the only ViT model adopting self-supervised training for this task, achieved accuracies of 82.2 and 84.6 when trained, respectively, on AudioSet-20K and AudioSet-2M (as reported in Table 2 of [21]). Since we did not have sufficient computational resources to pretrain RoMAE on AudioSet-2M, we compare its performance with that of the SSAST model pretrained on AudioSet-20K and Librispeech. We report the benchmark results in Table 9, showing that RoMAE performs better than the SSAST model in this context. Pre-training hyperparameters are reported in Table 8.
Table 9: Results for the ESC-50 benchmark for audio datasets.
Model Accuracy
SSAST (Librispeech) 80.0
RoMAE (Librispeech) 81.0
SSAST (AudioSet-20k) 82.2
RoMAE (AudioSet-20k) 83.4
A.5 Extrapolation
@@
Being a BERT-style model, RoMAE is not well-suited for extrapolation. During training the bidirectional encoder sees all tokens that lie inside the observed temporal window, therefore it never learns an inductive bias for causal ordering or forward progress in time, and struggles with out-of-distribution positions during inference. Recent work on causal Transformers, for example GPT-family models equipped with RoPE [52] or exponential relative embeddings [54, 53], shows that a strictly unidirectional attention pattern together with position embeddings that extrapolate to unseen indices can capture temporal trends far more effectively. We argue that despite this limitation, RoMAE has a place as a representation learning and interpolation framework, similarly to BERT in language.
Appendix B Proofs
@@
B.1 Reconstructing Absolute Position Using the [CLS] Token
@@
Proof: Consider a distance $r\in\mathbb{Q^{+}}\subset\mathbb{R}$, a query $\mathbf{q}=\boldsymbol{\psi}$ that is non-zero by assumption and a key corresponding to the [CLS] token as described in Definition B.1 such that $\mathbf{k}=\mathbf{R}^{r}\boldsymbol{\psi}$. Assume that the query is at position $j\in\mathbb{Q^{+}}$. We compute the dot product between rotated $\mathbf{q}$ and $\mathbf{k}$:
@@
Appendix C Full Experimental Details and Hyperparameters
C.1 Tiny Imagenet Experimental Setup
Table 10: Tiny ImageNet fine-tuning hyperparameters. Values that are constant across all runs are reported only once.
Hyperparameter RoMAE ([CLS]) RoMAE (no [CLS]) RoMAE (absolute)
Optimizer SGD
Momentum 0.9
Weight Decay 0.
Base LR $6\times 10^{-2}$ $6\times 10^{-2}$ $3\times 10^{-2}$
Batch size 512
Epochs 15
Gradient clip 1
Linear LR warmup steps 500
LR schedule Cosine
Dropout 0
Stochastic depth $\lambda$ 0
@@ C.1 Tiny Imagenet Experimental Setup
Precision FP16
Table 11: Tiny ImageNet unified pre-training hyperparameters.
Hyperparameter Pre-train
Optimizer AdamW
AdamW betas $\beta_{1}=0.9,\ \beta_{2}=0.95$
Weight Decay 0.05
Base LR $2\times 10^{-4}$
Batch size 1024
Epochs 200
Gradient clip 1
Linear LR warmup steps 2000
LR schedule Cosine
Dropout 0
Stochastic depth $\lambda$ 0
Label smoothing $c$ 1
Precision FP16
We present the Tiny ImageNet pre-training hyperparameters in Table 11 and fine-tuning hyperparameters in Table 10. All Tiny ImageNet pre-training runs use the same hyperparameters, while the fine-tuning runs use different learning rates. We find that FP16 precision works well here as opposed to our experiments on the ELAsTiCC2 dataset discussed in Section C.6.
When training, we normalize using the ImageNet mean and standard deviation. We also use RandAugment [13] with 2 operations and a magnitude of 9. While the model would likely benefit from more epochs during the pre-training and fine-tuning stages, we consider this sufficient for the purposes of an ablation study.
C.2 UEA Multivariate Time-series Experimental Setup

### v2

@@ 5.5 Interpolation
Table 7: Results for the interpolation experiments discussed in Section $5.$
@@
HeTVAE [35] $0.562\pm 0.022$}
RoMAE-tiny $\mathbf{0.467\pm 0.021}$
@@
RoMAE-tiny $\mathbf{0.0183\pm 0.007}$
We evaluate RoMAE on three interpolation tasks with increasing dimensionality and sampling irregularity. (i) Spiral: A 2D synthetic benchmark of 300 noisy Archimedean spirals as in Ref. [12]; (ii) Synthetic: The 50-step univariate task from Ref. [48] and (iii) PhysioNet: 48-hour ICU records containing 41 clinical variables [49]. For (i), each spiral is discretized into 75 evenly-spaced time steps. To create irregular time-series data, time points are randomly selected from the first half of each spiral, which are used for interpolation. For (ii), the interpolation task is between a random subsample including between 3 and 10 points per trajectory. For (iii), we follow the 50% masking protocol of [35]: half of the time, rows with at least one observation are hidden and must be reconstructed from context. We provide additional experimental details in the Appendix, respectively 19, D.8, and D.9. All experiments are conducted to ensure a direct comparison with the results of the Transformer [58], LatentODE [9], and ContiFormer [12] for (i), and HeTVAE for (ii) and (iii). Across scales, our RoMAE-tiny configuration consistently scores competitively with respect to benchmarked MSE/RMSE, as seen in Table. 7. We lastly remark on RoMAE’s ability to retain progressively higher frequency modes for interpolation with time series data in Appendix B.4.
@@ 6 Discussion
Table 7 shows that one tiny/small RoMAE model, pre-trained once with a generic masked-autoencoder objective and no task-specific architectural tuning, matches or surpasses specialised baselines across three increasingly difficult interpolation datasets. On the 2D Spiral benchmark, we improve by an order of magnitude over the previous best result by ContiFormer, which we attribute to MAE tubelet-masking enforcing long-range reasoning. For the 50-step Synthetic task we obtain $0.233\pm 0.007$ MSE, comparable to HeTVAE’s $0.223\pm 0.070$ but with markedly tighter standard deviation (five seeds). On the irregular, 41-channel PhysioNet-2012 ICU data we achieve $0.570\pm 0.014$ MSE, within half a standard deviation of HeTVAE’s heteroscedastic decoder. We have verified by retraining HeTVAE on the the PhysioNet interpolation task that it indeed excels on densely sampled variables, such as heart-rate, whereas RoMAE maintains more balanced interpolation across the sparsely observed channels; consequently, while the aggregate MSE over all the features is similar, RoMAE delivers a simpler, single-stage model without task-specific adaptations.
@@
Limitations: RoPE in RoMAE has some additional computational overhead if the positions are different with each forward pass, e.g., with any continuous irregular time-series. If positions stay constant, however, as in images, the overhead becomes negligible. For details on the additional compute incurred by continuous RoPE, see Appendix B.2. RoMAE is also not well suited for very long sequences, as it uses standard Attention which has $O(n^{2})$ memory complexity with regards to sequence length. Lastly, RoMAE’s ability to perform on extrapolation tasks is limited, as discussed in Sec. B.3.
@@ 7 Conclusion
Acknowledgments
RT and AS acknowledge funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. RT is partially supported by the Fondazione ICSC, Spoke 3 “Astrophysics and Cosmos Observations”, Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19)” - Next Generation EU (NGEU). We also acknowledge the use of computational resources provided by the Italian AREA Science Park supercomputing platform ORFEO in Trieste. GC is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Postdoctoral Fellowship Programme, SMASH co-funded under the grant agreement No. 101081355. The operation (SMASH project) is co-funded by the Republic of Slovenia and the European Union from the European Regional Development Fund.
Appendix A Model Details
A.1 Model Sizes
Table 8: All RoMAE model sizes.
Parameter tiny-shallow tiny small base
$d_{\text{model}}$ 180 180 432 720
$N_{\text{head}}$ 3 3 6 12
Depth 2 12 12 12
Dim. feed-forward 720 720 1728 2880
Num. parameters 0.782M 4.67M 26.9M 74.7M
A.2 Architectural, Normalization, and Regularization Details
@@
$\bar{a}_{i}=\frac{a_{i}}{\text{RMS}(\mathbf{a})}g_{i},\qquad\text{RMS}(\mathbf{a})=\sqrt{\epsilon+\frac{1}{d_{\text{model}}}\sum_{i=1}^{d_{\text{model}}}a_{i}^{2}}$ (3)
where $\mathbf{a}$ is a sequence of embeddings, $g_{i}$ is a learned parameter that rescales each $a_{i}\in\mathbf{a}$, and $\epsilon$ is a small value added for numerical stability. Notably, RMSNorm does not centre the input $\mathbf{a}$ like LayerNorm [2] does. Re-centring has been shown not to be necessary and removing it saves compute.
@@
Appendix B Additional Experiments and Discussion
B.1 Additional Absolute Position Reconstruction Results
@@
Here we provide an additional experiment showing how RoMAE is able to to reconstruct absolute position across a wide range of values. To conduct the experiment, we pass only one token into RoMAE, giving it a random position drawn from a uniform distribution $\mathcal{U}_{[a,\ b]}$, training the model to predict this position. We also pass in the [CLS] token. The experiment is conducted over two domains; (0, 100)(0,\ 100) and (0, 1000)(0,\ 1000). Hyperparameters and training details are discussed in Section D.5. After training, we evaluate how well the model performs across the range of values it was trained on. Results are plotted in Figure 2.
@@
Discussion: We find that the model is generally able to reconstruct all position values within the two domains tested, except when the position is close to the edges of the domain. This effect occurs already when the domain is relatively small and worsens as it becomes larger. We also find that the model performs better overall on the smaller domain. These results provide empirical support for Proposition 4.2, and show how the model is able to learn to reconstruct absolute position across a large domain. That the loss grows as the position nears the edges of the domain shows that the model does not find solutions that generalize to out-of-distribution positions. These results also indicate that it may be beneficial to rescale positions to be within a smaller range.
B.2 Compute Performance
Table 9: Relative speed of RoMAE when used with regular/irregular positions.
Positional Embedding Relative speed
Absolute (sin/cos) 1
RoPE (quantized) 0.98
RoPE (continuous) 0.87
We evaluate the performance of RoMAE when using different positional encoding methods, specifically: absolute sin/cos [58], RoPE with integer (quantized) positions, and RoPE with continuous positions. The workload we test on uses 2 positional dimensions for RoPE and 2D image-like inputs to the model. Therefore, this experiment is representative of what one would encounter when using data such as what we have in the Tiny ImageNet experiment (Section 5.2), or in the ELAsTiCC experiment (Section 5.4). The results, calculated on an NVIDIA 1650Ti GPU, are shown in Table 9.
Although the performance of regular quantized RoPE is not far from standard absolute positional embeddings, when switching to continuous RoPE the model is only 87% of the original speed. This is because we are unable to cache the RoPE frequencies between forward passes. With quantized position on the other hand, everything can be cached once before-hand and reused. While continuous RoPE incurs a notable performance penalty, it is not drastic. We note that other architectures specialized in irregular time-series also suffer from this issue, e.g., ContiFormer [11] is reported as being 6 times slower than the vanilla transformer. The performance of RoMAE could likely be improved through a more optimized RoPE implementation. Quantizing the positions could also address this issue in datasets where it is reasonable to do.
B.3 Extrapolation
@@
Being a BERT-style model, RoMAE is not well suited for extrapolation. During training the bidirectional encoder sees all tokens that lie inside the observed temporal window, therefore it never learns an inductive bias for causal ordering or forward progress in time, and struggles with out-of-distribution positions during inference. Recent work on causal Transformers, for example GPT-family models equipped with RoPE [52] or exponential relative embeddings [54, 53], shows that a strictly unidirectional attention pattern together with position embeddings that extrapolate to unseen indices can capture temporal trends far more effectively. We argue that despite this limitation, RoMAE has a place as a representation learning and interpolation framework, similarly to BERT in language.
B.4 Retaining Frequencies
Figure 3: Average MSE obtained from the interpolation task using RoMAE-tiny for time-series with a single varying frequency component. Left: MSE computed on the Fast Fourier Transform (FFT). Right: MSE in the time domain. We generate 200 time-series per individual frequency, with 50 observed noisy points and 50 masked (interpolated) points, thus a limiting frequency of 25 according to Nyquist-Shannon sampling theorem. Error bars show the standard deviation of the MSE obtained for each individual frequency.
Figure 4: Same as above but now for time-series with two frequency modes present in the signal. Left: MSE computed on the FFT. Right: MSE in the time domain. The time-series have 50 observed noisy points and 50 masked (interpolated) points.
Figure 5: Illustrative realisation from the evaluation of RoMAE on a the bi-frequency time series. Left: Interpolation in the time domain for a composite sinusoidal signal with base frequencies 1 and 5 Hz. Right: FFT of the ground truth and predicted waveform.
In this appendix we investigate ROMAE’s ability to retain high frequency modes in interpolation tasks. It is known, for example as shown in References [42, 64], that neural networks can exhibit a spectral bias, in that the networks preferentially learn low-frequency components before high-frequency details. Examining how RoMAE reconstructs patterns at different frequencies provides insight into whether the rotational encoding allows to capture fine-grained structure during interpolation, with implications for understanding the inductive biases introduced by this positional encoding scheme.
To empirically assess RoMAE’s ability to reconstruct signals at different frequencies, we designed a controlled toy dataset of noisy sine waves. Each time series is defined over $t\in[0,1]$ and generated as the sum of one or two frequency components (with equal probability), where integer frequencies are sampled uniformly from $f\in[0,24]$ Hz and Gaussian noise $\varepsilon\sim\mathcal{N}(0,0.01)$ is added. Each time-series has 100 data points, 50 of which are taken as input and 50 of which are masked for interpolation. We train RoMAE-tiny for 200 epochs on 10,000 examples. We then evaluate this model on (i) time-series with a single frequency mode as shown in Figure 3, (ii) time-series with two frequencies modes as shown in Figure 4. Using the 50 predicted (interpolated) points, we compute the MSE in both the time-domain and Fourier domain. We plot a sample prediction from the evaluation in Figure 5. This analysis was conducted with a mean signal to noise ratio (SNR) of 43.1 over the entire training set. We acknowledge that the distribution of the MSE will be affected by an increasing SNR, resulting in less sensitivity to higher frequency modes.
As expected, we observe a general degradation of the reconstruction for higher frequencies, with an approximately linear trend in error for frequency above 9Hz. For the case of two modes, we observe the same overall trend with a slight preference toward two higher frequency modes, as opposed to one low and one high mode. This is due to the sampling rate of the observations and the fact that the FFT for two higher frequencies has a uni-modal power spectrum, yielding slightly better reconstruction. Lastly, we have checked that the above observations are maintained for non-sinusoidal signals. We repeated the analysis using non-sinusoidal periodic functions, specifically a square wave and a cycloid. We observed similar behaviour for the retention of high frequencies as with the sinusoidal experiment.
Appendix C Proofs
@@
C.1 Reconstructing Absolute Position Using the [CLS] Token
@@
Proof: Consider a distance $r\in\mathbb{Q^{+}}\subset\mathbb{R}$, a query $\mathbf{q}=\boldsymbol{\psi}$ that is non-zero by assumption and a key corresponding to the [CLS] token as described in Definition C.1 such that $\mathbf{k}=\mathbf{R}^{r}\boldsymbol{\psi}$. Assume that the query is at position $j\in\mathbb{Q^{+}}$. We compute the dot product between rotated $\mathbf{q}$ and $\mathbf{k}$:
@@
Appendix D Full Experimental Details and Hyperparameters
D.1 Tiny Imagenet Experimental Setup
Table 10: Tiny ImageNet unified pre-training and fine-tuning hyperparameters.
Hyperparameter Pre-Training Fine-Tuning
Optimizer AdamW AdamW
AdamW betas $\beta_{1}=0.9,\ \beta_{2}=0.95$ $\beta_{1}=0.9,\ \beta_{2}=0.999$
Weight Decay 0.05 0.05
Base LR $2\times 10^{-4}$ $1\times 10^{-3}$
Batch size 1024 1024
Epochs 200 15
Gradient clip 1 1
Linear LR warmup steps 2000 500
LR schedule Cosine Cosine
Dropout 0 0
Stochastic depth $\lambda$ 0 0
@@ D.1 Tiny Imagenet Experimental Setup
Precision FP32 FP32
We present the unified Tiny ImageNet pre-training and fine-tuning hyperparameters in Table 10. All Tiny ImageNet pre-training and fine-tuning runs use the same hyperparameters. Although our final results use FP32, we also tested mixed-precision training and found that FP16 precision works with Tiny-ImageNet as opposed to our experiences on the ELAsTiCC dataset discussed in Section D.6.
When training, we normalize using the ImageNet mean and standard deviation. Each patch is individually normalized when calculating loss as is done in MAE [23]. During fine-tuning, we also use RandAugment [13] with 2 operations and a magnitude of 9. While the model would likely benefit from more epochs during the pre-training and fine-tuning stages, we consider this sufficient for the purposes of an ablation study.
D.2 Audio Experimental Setup
In Table 11 we present the unified pre-training and fine-tuning hyperparameters for the audio representation learning and classification experiments, discussed in Section 5.3, in the main text.
When training, both on Audioset [20] and Librispeech [36], we normalize the data using the mean and standard deviation estimated on the entire set of spectrograms. Each patch is individually normalized before calculating the loss as is done in MAE [23].
Table 11: Audio Experiment unified pre-training and fine-tuning hyperparameters.
Hyperparameter Pre-Training Fine-Tuning
Optimizer AdamW AdamW
AdamW betas $\beta_{1}=0.9,\ \beta_{2}=0.95$ $\beta_{1}=0.9,\ \beta_{2}=0.99$
Weight Decay 0.05 0.02
Base LR $5\times 10^{-4}$ $1\times 10^{-3}$
Batch size 64 48
Epochs 150 50
Gradient clip 1 1
Linear LR warmup steps 1000 50
LR schedule Cosine Cosine
Dropout 0 0
Stochastic depth $\lambda$ 0 0
Label smoothing $c$ 0.8
Precision FP32 FP32
D.3 UEA Multivariate Time-series Experimental Setup